OpenStudy (anonymous):

Which is the exact value of the expression the square root of 32. − the square root of 50. + the square root of 128.? 2the square root of 7. 7the square root of 2. 22the square root of 2. 2the square root of 55.

OpenStudy (anonymous):

@Omar_Elboredy could you help with this one also?

OpenStudy (anonymous):

\[\sqrt{32}-\sqrt{50} + \sqrt{128} = 7\sqrt{2}\] which will be the second question (B)